# Local logistic regression for dimension reduction in classification

### v1

@@
Abstract. Sufficient dimension reduction has received much interest over the past 30 years. Most existing approaches focus on statistical models linking the response to the covariate through a regression equation, and as such are not adapted to binary classification problems. We address the question of dimension reduction for binary classification by fitting a localized nearest-neighbor logistic model with $\ell_{1}$-penalty in order to estimate the gradient of the conditional probability of interest. Our theoretical analysis shows that the pointwise convergence rate of the gradient estimator is optimal under very mild conditions. The dimension reduction subspace is estimated using an outer product of such gradient estimates at several points in the covariate space. Our implementation uses cross-validation on the misclassification rate to estimate the dimension of this subspace. We find that the proposed approach outperforms existing competitors in synthetic and real data applications.
@@ 1 Introduction
A number of statistical problems require predicting whether an event of interest $A$ happens given auxiliary information in the form of a random covariate $X\in\mathbb{R}^{p}$. In finance, for example, it is of interest to predict whether a given customer, characterized by information such as age, occupation and net income, is going to default on a loan. In hydrology, one may have to predict whether flood defenses are going to be breached during a storm based on weather variables and local topographical information. Marketing studies often try to assess customer churn within slices of a customer database, which amounts to predicting whether a customer is going to stop using a product given their past behavior and responses to marketing communications. In each of those problems, a natural approach is to set $Y=\mathbbm{1}_{A}$, the indicator function of $A$, and to model the conditional probability $\mathbb{P}(Y=1\,|\,X)$. A first, naive solution is then to model this conditional probability as $\mathbb{P}(Y=1\,|\,X)=\exp(\beta^{T}X)/(1+\exp(\beta^{T}X))$ and to estimate the direction $\beta\in\mathbb{R}^{p}$ with logistic regression. This is a well-understood methodology, part of generalized linear modelling, for which an excellent reference is Agresti (2013). Logistic regression, and its alternatives such as probit regression or log-log regression, is sensitive to misspecification like any other parametric technique. Moreover, it relies on all the information brought by $X$ about $Y$ to lie in a single projection of $X$, meaning that up to rotation within the covariate space, only a single component of $X$ should contain information about the event of interest. These strong restrictions have motivated the construction and study of several nonparametric and semiparametric regression methods over the past 30 years.
@@
Many such methods are partly built on a class of techniques aimed at estimating so-called sufficient dimension reduction subspaces. A dimension reduction subspace $E$ for the distribution of a random pair $(X,Y)\in\mathbb{R}^{p}\times\mathbb{R}$ is any linear subspace of $\mathbb{R}^{p}$ spanned by the ($p-$dimensional) columns of a matrix $B\in\mathbb{R}^{p\times d}$ with $p$ rows and $d$ columns such that $Y$ and $X$ are conditionally independent in the sense of Dawid (1979) given the vector $B^{T}X\in\mathbb{R}^{d}$; the notion of conditional independence, which is linked to fundamental statistical notions such as sufficiency and ancillarity, is also discussed from a different viewpoint in Basu and Pereira (1983). Of course, dimension reduction subspaces always exist and are not unique in general, but the central subspace, that is, the intersection of all dimension reduction subspaces, is identifiable and is in many cases a dimension reduction subspace as well (Cook, 1994, 1996; Portier and Delyon, 2013). An interesting case occurs when the central subspace has dimension $d=1$, which is the so-called single-index regression situation (Hall, 1989) where the distribution of $Y$ given $X$ is an unknown function of a single, unknown projection $\beta^{T}X$ to be estimated. This model extends traditional parametric regression models by allowing the link function $g$ satisfying $\mathbb{E}(Y|X)=g(\beta^{T}X)$ to be unknown. More generally, whenever a matrix $B$ forms a dimension reduction subspace, $Y$ can in fact be written as a measurable function of $B^{T}X$ and $\varepsilon$, where $\varepsilon$ is independent from $X$; this makes the model $Y=g(B^{T}X)+\varepsilon$, with $g:\mathbb{R}^{d}\to[0,1]$ unknown, a particularly appealing model for dimension reduction.
Among dimension reduction methods, Sliced Inverse Regression (SIR), pioneered by Li (1991), seeks to estimate the central subspace by recovering the conditional expectation $\mathbb{E}(X|Y)$ from a family of moments $\mathbb{E}(\psi(Y)X)$ depending on a class of (indicator) functions $\psi$ encoding whether $Y$ belongs to a particular slice of the data. Several extensions have been proposed, including Sliced Average Variance Estimation (SAVE) by Cook and Weisberg (1991). These methods rely on the strong assumption that $X$ has an elliptically symmetric distribution (Cook and Weisberg, 1991). Directional Regression (DR) approaches as in Li and Wang (2007) make assumptions that are weaker but related to elliptical symmetry, and Likelihood Acquired Directions (LAD) as in Cook and Forzani (2009) assume that $X$ given $Y$ is Gaussian. Differently from these techniques, the Principal Hessian Direction (PHD) approach developed by e.g. Li (1992) for Gaussian multivariate data, and further investigated by Cook (1998) under conditions akin to elliptical symmetry again, directly work on the conditional expectation $\mathbb{E}(Y|X)$ in order to recover the central subspace. Li (2007) modifies these methods to obtain sparse (with respect to the covariate index) dimension reduction subspaces. Adaptations of the SIR, SAVE, and PHD methods to rare event classification were studied by Aghbalou et al. (2024).
@@
Gradient-based estimation methods do not assume elliptical symmetry and are based on the estimation of the gradient of a regression function in order to recover vectors in the space of interest. Gradient estimation methods have been introduced by Härdle and Stoker (1989) and Powell et al. (1989); the works by Klein and Spady (1993) and Huh and Park (2002) focus specifically on binary regression but only allows for a one-dimensional dimension reduction subspace. These ideas have been extended in various ways using local linear estimation, see for example Hristache et al. (2001), which improves the average derivative estimator by focusing on the hyperplane orthogonal to the target direction in a single-index model, the Minimum Average Variance Estimation (MAVE) procedure of Xia et al. (2002) that simultaneously estimates the conditional variance of the observations given directions in the data and minimizes this quantity with respect to these directions, the method of Dalalyan et al. (2008) combining local linear estimation with an $\ell_{\infty}-$type Principal Component Analysis technique, and nearest-neighbor local linear estimators combined with a LASSO penalization devised by Ausset et al. (2021) in order to handle the high-dimensional setting. Local linear estimation, for which a comprehensive reference is Fan and Gijbels (1996), provides an efficient way to estimate the derivatives of the regression function and as such, it has an essential place in many dimension reduction methods such as those we have just highlighted. Somewhat tangentially, Delecroix et al. (2006) also build estimators based on minimizing general loss functions. With the exception of the work by Klein and Spady (1993) and Delecroix et al. (2006), the gradient-based methods outlined here are designed for dimension reduction in location regression models of the form $Y=g(\beta^{T}X)+\varepsilon$, which excludes the binary regression and classification frameworks; however, Klein and Spady (1993) and Delecroix et al. (2006) only handle single-index classification.
A more recent line of work has focused on using powerful modern tools of mathematical statistics such as Reproducing Kernel Hilbert Spaces (RKHS). The use of RKHS in dimension reduction appears to have been pioneered by Fukumizu et al. (2004), who use the covariance operator to characterize the central subspace when the response and covariate are finite-dimensional. Later, Hsing and Ren (2009) paired the use of RKHS with SIR to handle dimension reduction when the covariate is a stochastic process. Subsequent work by Li et al. (2011) unified the construction of linear and nonlinear dimension reduction subspaces again using support vector machine approaches that can handle binary regression models; this technique is closely linked to the Principal Optimal Transport Direction (POTD) method investigated in Meng et al. (2020). The recent solution put forward by Kang and Shin (2022), and building on earlier work by Mukherjee and Wu (2006), makes use of local estimation techniques based on an RKHS representation theorem to fit the function as well as its gradient. Their theory requires the true functions and gradient to lie within the RKHS space of interest, and to specify the RKHS kernel correctly. Computationally speaking, their optimization procedure relies on a risk function that is a second-order $U$-statistic involving $n^{2}$ terms. Other machine learning techniques, such as ensemble learning, have also been paired with local estimation (see for example Cai et al., 2023) although they tend to rely on local least squares and thus not on loss functions better suited to binary classification.
This article proposes a new dimension reduction framework tailored to classification, handling dimension reduction subspaces not limited to a single projection, and able to perform variable selection while dealing with the curse of dimensionality. Assume that $Y$ satisfies
$\pi(X):=\mathbb{P}(Y=1|X)=g(\beta^{T}X)$
@@
for some $\beta\in\mathbb{R}^{p\times d}$, $d\leq p$ and a measurable function $g:\mathbb{R}^{d}\to[0,1]$. The function $g$, treated as a nuisance parameter, and the matrix $\beta$ are both unknown. If we had $d=1$ and $g(t)=\operatorname{expit}(t):=\exp(t)/(1+\exp(t))$, then we would have $\operatorname{logit}(\pi(X)):=\log(\pi(X)/(1-\pi(X)))=\beta^{T}X$, and $\beta\in\mathbb{R}^{p}$ would be exactly the gradient of $\ell:=\operatorname{logit}(\pi)$. Our main idea is that, if the conditional probability $\pi$ is sufficiently regular at a given point $x$ then $\ell(X)$ can be approximated by a linear function of $X-x$ in a neighborhood of $x$, whose intercept $a=a(x)$ is just $\ell(x)$ and whose gradient $b=b(x)=\nabla\ell(x)$ belongs to the vector space $\operatorname{span}(\beta)$ that is spanned by the columns of $\beta$. We then construct, at each point $x$, a nearest-neighbor, penalized, local logistic (and hence convex) loss function depending on two arguments $a\in\mathbb{R}$ and $b\in\mathbb{R}^{p}$. The weak convergence of the estimator follows, under very mild conditions, from a new general empirical process theory result on a family of nearest-neighbor estimators that is of independent interest. The gradient estimates at several points are then aggregated using the standard outer product of gradients of e.g. Section 3.1 in Xia et al. (2002), and $\operatorname{span}(\beta)$ is recovered by finding the eigenspaces of this matrix corresponding to its $d$ largest eigenvalues. Using nearest neighbors (see Biau and Devroye (2015) for a recent textbook), instead of the Nadaraya-Watson weights as in Fan et al. (1995), ensures an adaptive bandwidth choice while keeping computational time reasonable thanks to algorithms such as $k-d$ tree search. Our use of a penalized loss function ensures sparsity of the estimated vectors in a similar spirit as in Li (2007) but without requiring a first step non-sparse estimation and, unlike earlier local generalized linear techniques such as that of Fan et al. (1995), makes it possible to deal with potentially large dimensions (of the order of several dozens when the sample size is $n=1000$). While most of the existing dimension reduction methods select the number $d$ of components in the reduction subspace by using rank testing procedures (Bura and Yang, 2011; Portier and Delyon, 2014), we here propose to rely on the underlying classification context by developing a simple cross-validation approach comparing the different sets of ordered eigenvectors given by the outer product of gradients. Surprisingly, this is helpful in improving the final accuracy of the classification method even when the true dimension of the reduction subspace is known.
@@
The paper is organized as follows. In Section 2.1, we describe our statistical framework in detail and we then construct our nearest-neighbor, penalized local logistic loss function, whose maximization gives rise to the proposed gradient estimator. In Section 2.2, we establish the pointwise weak convergence of this estimator and in Section 2.3 we employ an outer-product of gradients to estimate the central subspace. In Section 3, we discuss several practical aspects of the proposed algorithm, we construct our classification method, and we compare our approach to several competitors on synthetic and real data. The mathematical proofs and additional finite-sample results are postponed to an online Supplementary Material document.
@@ 2.1 Background
Let $Y\in\{0,1\}$ be a binary response variable, with random covariate $X\in\mathbb{R}^{p}$. The data at hand is supposed to be made of independent copies $(Y_{i},X_{i})$ of the random pair $(Y,X)$, for $1\leq i\leq n$. Suppose that for an $x$ in the support $S_{X}$ of $X$ (assumed to have nonempty interior),
$\pi(x)=\mathbb{P}(Y=1|X=x)={g(\beta^{T}x)},$
for some unknown $\beta\in\mathbb{R}^{p\times d}$, $d\leq p$ and an unknown measurable function $g:\mathbb{R}^{d}\to[0,1]$. The Kullback-Leibler divergence of the Bernoulli distribution with parameter $\pi\in(0,1)$ from the Bernoulli distribution with parameter $q\in(0,1)$ is
@@
For $d=1$ and $g_{0}=\operatorname{expit}$, this is the logistic regression estimator, with the empirical integrated cross-entropy loss being just the negative conditional log-likelihood in the logistic model.
When $g$ is unknown, the logistic regression estimator is not guaranteed to be a consistent estimator of $\beta$. However, if $\pi$ is differentiable at $x$, then the gradient $\nabla\pi(x)$ of $\pi$ at $x$ satisfies
$\nabla\pi(x)=\beta\nabla g(\beta^{T}x).$
As such, $\nabla\pi(x)\in\operatorname{span}(\beta)$. If moreover $10<\pi(x)<$ 1, then likewise
@@
to be minimized at $(a_{\varepsilon}(x),b_{\varepsilon}(x))\approx(\ell(x),\nabla\ell(x))$ as $\varepsilon\downarrow 0$, where $B(x,\varepsilon)$ is the closed Euclidean ball with center $x$ and radius $\varepsilon$ in $\mathbb{R}^{p}$. This shall be the cornerstone of our methodology.
@@ 2.2 Nearest-neighbor penalized local logistic estimator
The localization induced by conditioning upon covariate values belonging to $B(x,\varepsilon)$ is reproduced empirically using a nearest-neighbor procedure. Fix $x\in\mathbb{R}^{p}$ and let $N_{k}(x)\subset\{1,\ldots,n\}$ be the set gathering the indices of the $k$-nearest neighbors $X_{i}$ of the point $x$; we shall assume in the theory below that the distribution of $X$ has a density w.r.t. Lebesgue measure, so that ties will not happen with probability 1 and $N_{k}(x)$ is well-defined. The empirical counterpart of the local integrated cross-entropy is then
@@
$L_{n}(a,b)=\sum_{i\in N_{k}(x)}Y_{i}\log(\operatorname{expit}(a+b^{T}(X_{i}-x)))+(1-Y_{i})\log(1-\operatorname{expit}(a+b^{T}(X_{i}-x))).$
@@
The asymptotic behavior of $\overline{L}_{n}(a,b)$ is much easier to study than that of $L_{n}(a,b)$, since it is a sum of independent and identically distributed random variables; like $L_{n}(a,b)$, it defines a concave objective function and therefore one should expect that (up to technical details) the asymptotic behavior of $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$ will follow from the pointwise convergence of $\overline{L}_{n}(a,b)$. The key result in order to make this intuition rigorous is of independent interest and is the first main result of this paper. Here and throughout $\|\cdot\|_{2}$ denotes the standard Euclidean norm.
Theorem 1 (Central limit theorem for nearest-neighbor estimators).
Suppose that (A1) is fulfilled. Assume that $k:=k_{n}\to\infty$ is such that $k/n\to 0$. Let $\Psi_{n}:\{0,1\}\times\mathbb{R}^{p}\to\mathbb{R}^{q}$ be a sequence of measurable vector-valued functions and suppose that there is a positive integer $n_{0}$ such that
$\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\max(\|\Psi_{n}(0,z)\|_{2},\|\Psi_{n}(1,z)\|_{2})<\infty.$
  1. (i)
    Then $Z_{n}(\widehat{\tau}_{n,k}(x))=O_{\mathbb{P}}(1)$.
  2. (ii)
    then
Theorem 1(ii) is obtained by, first, showing that the empirical process $(Z_{n}(\tau))$ is tight and converges weakly to a continuous Gaussian process within a suitable functional space when the radius $\tau$ is restricted to being of similar order as the nearest neighbor radius $\tau_{n,k}(x)$, and then by combining that with the fact that $\widehat{\tau}_{n,k}(x)/\tau_{n,k}(x)\to 1$ in probability, see Lemma 2.
@@
where $0_{p}$ is the zero vector in $\mathbb{R}^{p}$ and $I_{p}$ denotes the identity matrix of order $p$. Our first main result provides the limiting distribution of the pair $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$. In this result, assumption (A2) ensures that the gradient of $\ell=\operatorname{logit}\pi$ is well-defined, with the second order derivatives of $\pi$ and hence of $\ell$ coming into play when evaluating the rate of convergence of the bias term incurred when localizing. Denote by $\Delta\pi(x)$ the Laplacian of $\pi$ at $x$.
@@
The conditions on $k$ and $\lambda$ in Theorem 2 are equivalent to assuming that $\tau_{n,k}(x)\sqrt{k}\to\infty$, $\tau_{n,k}^{2}(x)\sqrt{k}$ is bounded and $\lambda/(\tau_{n,k}(x)\sqrt{k})$ converges to the finite constant $(cf_{X}(x)V_{p})^{1/p}$. Condition $\tau_{n,k}(x)\sqrt{k}\to\infty$ is necessary in order to be able to write a Taylor expansion of the penalized component in the loss function (although it is, strictly speaking, not needed for the analysis of the non-penalized local logistic maximum likelihood function, see Lemma 5). In pointwise results on local linear kernel quasi-maximum likelihood estimation with one-dimensional covariates, Theorem 1a in Fan et al. (1995) requires $\sqrt{nh^{3}}=h\sqrt{nh}\to\infty$, where $h$ is the kernel bandwidth; note that $h$ and $\tau_{n,k}(x)$ play the same role and that, for $p=1$, the kernel regression analogue of $k$ is a quantity proportional to $nh$, so that conditions $h\sqrt{nh}\to\infty$ and $\tau_{n,k}(x)\sqrt{k}\to\infty$ are indeed analogous. Of course, condition $\tau_{n,k}(x)\sqrt{k}\to\infty$ is automatically satisfied if $\tau_{n,k}^{2}(x)\sqrt{k}$ converges to a finite positive limit, namely, when the nonparametric bias-variance tradeoff is achieved. We also highlight that condition $k/n\to 0$, which is standard in nearest-neighbor estimation, follows from assuming that $k\to\infty$ and $\tau_{n,k}^{2}(x)\sqrt{k}$ is bounded.

### v2

# Logistic lasso regression with nearest neighbors for gradient-based dimension reduction

@@
Abstract. This paper investigates a new approach to estimate the gradient of the conditional probability given the covariates in the binary classification framework. The proposed approach consists in fitting a localized nearest-neighbor logistic model with $\ell_{1}$-penalty in order to cope with possibly high-dimensional covariates. Our theoretical analysis shows that the pointwise convergence rate of the gradient estimator is optimal under very mild conditions. Moreover, using an outer product of such gradient estimates at several points in the covariate space, we establish the rate of convergence for estimating the so-called central subspace, a well-known object allowing to carry out dimension reduction within the covariate space. Our implementation uses cross-validation on the misclassification rate to estimate the dimension of this subspace. We find that the proposed approach outperforms existing competitors in synthetic and real data applications.
@@ 1 Introduction
A number of statistical problems require predicting whether an event of interest $A$ happens given auxiliary information in the form of a random covariate $X\in\mathbb{R}^{p}$. Starting with logistic regression, many methods have been developed in order to solve this so-called binary classification problem. With the now routine collection of high-dimensional data, many such methods are partly built on a class of techniques aimed at estimating so-called sufficient dimension reduction subspaces. A dimension reduction subspace $E$ for the distribution of a random pair $(X,Y)\in\mathbb{R}^{p}\times\mathbb{R}$ is any linear subspace of $\mathbb{R}^{p}$ spanned by the ($p-$dimensional) columns of a matrix $B\in\mathbb{R}^{p\times d}$ with $p$ rows and $d$ columns such that $Y$ and $X$ are conditionally independent (Dawid, 1979) given the vector $B^{T}X\in\mathbb{R}^{d}$; the notion of conditional independence, which is linked to fundamental statistical notions such as sufficiency and ancillarity, is also discussed from a different viewpoint in Basu and Pereira (1983). Of course, dimension reduction subspaces always exist and are not unique in general, but the central subspace, that is, the intersection of all dimension reduction subspaces, is identifiable and is in many cases a dimension reduction subspace as well (Cook, 1994, 1996; Portier and Delyon, 2013). When $d=1$, which is the so-called single-index regression situation (Hall, 1989), the distribution of $Y$ given $X$ is an unknown function of a single, unknown projection $\beta^{T}X$ to be estimated. This model extends traditional parametric regression models by allowing the link function $g$ satisfying $\mathbb{E}(Y|X)=g(\beta^{T}X)$ to be unknown. More generally, whenever a matrix $B$ forms a dimension reduction subspace, $Y$ can in fact be written as a measurable function of $B^{T}X$ and $\varepsilon$, where $\varepsilon$ is independent from $X$; this makes the model $Y=g(B^{T}X)+\varepsilon$, with $g:\mathbb{R}^{d}\to[0,1]$ unknown, particularly appealing for dimension reduction.
@@
Among the existing classes of dimension reduction methods, gradient-based estimation methods target the gradient of a regression function in order to recover vectors in the space of interest. They have been introduced by Härdle and Stoker (1989) and Powell et al. (1989) (see also gasser+m:1984 for earlier results on gradient estimation); the works by Klein and Spady (1993) and Huh and Park (2002) focus specifically on binary regression but only allows for a one-dimensional dimension reduction subspace. These ideas have been extended in various ways using local linear estimation (Fan and Gijbels, 1996), see for example Hristache et al. (2001), which improves the average derivative estimator by focusing on the hyperplane orthogonal to the target direction in a single-index model, the Minimum Average Variance Estimation (MAVE) procedure of Xia et al. (2002) that simultaneously estimates the conditional variance of the observations given directions in the data and minimizes this quantity with respect to these directions, the method of Dalalyan et al. (2008) combining local linear estimation with an $\ell_{\infty}-$type Principal Component Analysis technique, and nearest-neighbor local linear estimators combined with a LASSO penalization devised by Ausset et al. (2021) in order to handle the high-dimensional setting. Local linear estimation provides an efficient way to estimate the derivatives of the regression function, and as such, it has an essential place in many dimension reduction methods such as those we have just highlighted. Somewhat tangentially, Delecroix et al. (2006) also build estimators based on minimizing general loss functions. With the exception of the work by Klein and Spady (1993) and Delecroix et al. (2006), the gradient-based methods outlined here are designed for dimension reduction in location regression models of the form $Y=g(\beta^{T}X)+\varepsilon$, which excludes the binary regression and classification frameworks; however, Klein and Spady (1993) and Delecroix et al. (2006) only handle single-index classification.
@@
This article proposes a new dimension reduction framework tailored to classification, handling dimension reduction subspaces not limited to a single projection, and able to perform variable selection while dealing with the curse of dimensionality. Assume that $Y$ satisfies $\pi(X):=\mathbb{P}(Y=1|X)=g(\beta^{T}X),$ for some $\beta\in\mathbb{R}^{p\times d}$, $d\leq p$ and a measurable function $g:\mathbb{R}^{d}\to[0,1]$. The function $g$, treated as a nuisance parameter, and the matrix $\beta$ are both unknown. If we had $d=1$ and $g(t)=\operatorname{expit}(t):=\exp(t)/(1+\exp(t))$, then we would have $\operatorname{logit}(\pi(X)):=\log(\pi(X)/(1-\pi(X)))=\beta^{T}X$, and $\beta\in\mathbb{R}^{p}$ would be exactly the gradient of $\ell:=\operatorname{logit}(\pi)$. Our main idea is that, if the conditional probability $\pi$ is sufficiently regular at a given point $x$ then $\ell(X)$ can be approximated by a linear function of $X-x$ in a neighborhood of $x$, whose intercept $a=a(x)$ is just $\ell(x)$ and whose gradient $b=b(x)=\nabla\ell(x)$ belongs to the vector space $\operatorname{span}(\beta)$ that is spanned by the columns of $\beta$. We then construct, at each point $x$, a nearest-neighbor, penalized, local logistic (and hence convex) loss function depending on two arguments $a\in\mathbb{R}$ and $b\in\mathbb{R}^{p}$. The weak convergence of the estimator of $(\ell(x),\nabla\ell(x))$ hence constructed follows, under very mild conditions, from a new general empirical process theory result on a family of nearest-neighbor estimators that is of independent interest. The obtained estimation rates for $\ell(x)$ and $\nabla\ell(x)$ achieve the minimax optimal rates of convergence as described in Stone (1982). In particular, our theoretical results compare favorably to the ones established in Kang and Shin (2022), building on earlier work by Mukherjee and Wu (2006), where the gradient is estimated using a Reproducing Kernel Hilbert Spaces (RKHS) technique (see below our Corollary 2 for a precise comparison). The gradient estimates at several points are then aggregated using the standard outer product of gradients of e.g. Section 3.1 in Xia et al. (2002), and $\operatorname{span}(\beta)$ is recovered by finding the eigenspaces of this matrix corresponding to its $d$ largest eigenvalues. Using the nearest neighbor method, as introduced in fix1989discriminatory and studied in textbooks such as devroye2013probabilistic and Biau and Devroye (2015), instead of the Nadaraya-Watson weights as in Fan et al. (1995), ensures an adaptive bandwidth choice while keeping computational time reasonable thanks to algorithms such as $k-d$ tree search. Our penalization ensures sparsity of the estimated vectors in a similar spirit as in Li (2007) but without requiring a first step non-sparse estimation and, unlike earlier local generalized linear techniques such as that of Fan et al. (1995), makes it possible to deal with large dimensions (of the order of several dozens when the sample size is $n=1000$). While most existing dimension reduction methods select the number $d$ of components in the reduction subspace by using rank testing procedures (Bura and Yang, 2011; Portier and Delyon, 2014), we here propose to rely on the underlying classification context by developing a simple cross-validation approach comparing the different sets of ordered eigenvectors given by the outer product of gradients. Surprisingly, this is helpful in improving the final accuracy of the classification method even when the true dimension of the reduction subspace is known.
A somewhat orthogonal approach to dimension reduction, called Sliced Inverse Regression (SIR), pioneered by Li (1991) and Cook and Weisberg (1991), seeks to estimate the central subspace by recovering the conditional expectations $\mathbb{E}(X|Y)$ or $\mathbb{E}(XX^{T}|Y)$ from a family of moments. The main drawback of such an approach, in contrast to the one proposed here, is that the covariates usually need to satisfy the so-called linearity condition, e.g., by having an elliptically symmetric distribution. We refer to li2018sufficient for a textbook treatment of inverse regression methods. Note also another line of work relying on RKHS, pioneered by Fukumizu et al. (2004), who use the covariance operator to characterize the central subspace and, by doing so, avoiding the reliance on the linearity condition. Using a support vector machine approach, Li et al. (2011) unified the construction of linear and nonlinear dimension reduction subspaces that can handle binary regression models; this technique is closely linked to the Principal Optimal Transport Direction (POTD) method investigated in Meng et al. (2020).
@@
The paper is organized as follows. In Section 2.1, we describe our statistical framework and we construct our nearest-neighbor, penalized local logistic loss function, whose maximization gives rise to the proposed gradient estimator. In Section 2.2, we establish the pointwise weak convergence of this estimator and in Section 2.3 we employ an outer-product of gradients to estimate the central subspace. In Section 3, we discuss the practical aspects of the proposed algorithm, we construct our classification method, and we compare our approach to several competitors on synthetic and real data. Mathematical proofs and additional finite-sample results are postponed to the online Supplementary Material document.
@@ 2.1 Background
Let $Y\in\{0,1\}$ be a binary response variable, with random covariate $X\in\mathbb{R}^{p}$. The data is assumed to be made of independent copies $(Y_{i},X_{i})$ of the random pair $(Y,X)$, for $1\leq i\leq n$. Suppose that for an $x$ in the support $S_{X}$ of $X$ (assumed to have nonempty interior), $\pi(x)=\mathbb{P}(Y=1|X=x)=g(\beta^{T}x),$ for some unknown $\beta\in\mathbb{R}^{p\times d}$, $d\leq p$ and an unknown measurable function $g:\mathbb{R}^{d}\to[0,1]$. The Kullback-Leibler divergence of the Bernoulli distribution with parameter $\pi\in(0,1)$ from the Bernoulli distribution with parameter $q\in(0,1)$ is
@@
For $d=1$ and $g_{0}=\operatorname{expit}$, this is the logistic regression estimator, and the empirical integrated cross-entropy loss is the negative conditional log-likelihood in the logistic model.
When $g$ is unknown, the logistic regression estimator is not guaranteed to be a consistent estimator of $\beta$. However, if $\pi$ is differentiable at $x$, then the gradient $\nabla\pi(x)$ of $\pi$ at $x$ satisfies $\nabla\pi(x)=\beta\nabla g(\beta^{T}x).$ As such, $\nabla\pi(x)\in\operatorname{span}(\beta)$. If moreover $10<\pi(x)<$ 1, then likewise
@@
to be minimized at $(a_{\varepsilon}(x),b_{\varepsilon}(x))\approx(\ell(x),\nabla\ell(x))$ as $\varepsilon\downarrow 0$, where $B(x,\varepsilon)$ is the closed Euclidean ball with center $x$ and radius $\varepsilon$ in $\mathbb{R}^{p}$. This is the cornerstone of our methodology.
@@ 2.2 Nearest-neighbor penalized local logistic estimator
The localization induced by conditioning upon covariate values belonging to $B(x,\varepsilon)$ is reproduced empirically using the well known nearest-neighbor procedure. Fix $x\in\mathbb{R}^{p}$ and let $N_{k}(x)\subset\{1,\ldots,n\}$ be the set gathering the indices of the $k$-nearest neighbors $X_{i}$ of the point $x$; we shall assume in the theory below that the distribution of $X$ has a density w.r.t. Lebesgue measure, so that ties will not happen with probability 1 and $N_{k}(x)$ is well-defined. The empirical counterpart of the local integrated cross-entropy is then
@@
$L_{n}(a,b)$ $=\sum_{i\in N_{k}(x)}Y_{i}\log(\operatorname{expit}(a+b^{T}(X_{i}-x)))+(1-Y_{i})\log(1-\operatorname{expit}(a+b^{T}(X_{i}-x)))$
$=\sum_{i\in N_{k}(x)}Y_{i}(a+b^{T}(X_{i}-x))-\log(1+\exp(a+b^{T}(X_{i}-x))).$ (2)
@@
The asymptotic behavior of $\overline{L}_{n}(a,b)$ is much easier to study than that of $L_{n}(a,b)$, since it is a sum of independent and identically distributed random variables; like $L_{n}(a,b)$, it defines a concave objective function and therefore one should expect that (up to technical details) the asymptotic behavior of $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$ will follow from the pointwise convergence of $\overline{L}_{n}(a,b)$. The key result in order to make this intuition rigorous, which is Theorem 4, is deferred to the Supplementary material.
@@
where $0_{p}$ is the zero vector in $\mathbb{R}^{p}$ and $I_{p}$ denotes the identity matrix of order $p$. Our first main result provides the limiting distribution of the pair $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$. In this result, assumption (A2) ensures that the gradient of $\ell=\operatorname{logit}\pi$ is well-defined, with the second order derivatives of $\pi$ and hence of $\ell$ coming into play when evaluating the bias term incurred when localizing. Denote by $\Delta\pi(x)$ the Laplacian of $\pi$ at $x$ and by $\|\cdot\|_{2}$ the Euclidean norm.
@@
The conditions on $k$ and $\lambda$ in Theorem 1 are equivalent to assuming that $\tau_{n,k}(x)\sqrt{k}\to\infty$, $\tau_{n,k}^{2}(x)\sqrt{k}$ is bounded and $\lambda/(\tau_{n,k}(x)\sqrt{k})$ converges to the finite constant $(cf_{X}(x)V_{p})^{1/p}$. Condition $\tau_{n,k}(x)\sqrt{k}\to\infty$ is necessary in order to be able to write a Taylor expansion of the penalized component in the loss function (this is not needed for the analysis of the non-penalized local logistic maximum likelihood function, see Lemma 5). In pointwise results on local linear kernel quasi-maximum likelihood estimation with one-dimensional covariates, Theorem 1a in Fan et al. (1995) requires $\sqrt{nh^{3}}=h\sqrt{nh}\to\infty$, where $h$ is the kernel bandwidth; note that $h$ and $\tau_{n,k}(x)$ play the same role and that, for $p=1$, the kernel regression analogue of $k$ is a quantity proportional to $nh$, so that conditions $h\sqrt{nh}\to\infty$ and $\tau_{n,k}(x)\sqrt{k}\to\infty$ are indeed analogous, and constitute the assumptions required in order to ensure consistency of the local linear estimators of $\ell$ and its gradient. Of course, condition $\tau_{n,k}(x)\sqrt{k}\to\infty$ is automatically satisfied if $\tau_{n,k}^{2}(x)\sqrt{k}$ converges to a finite positive limit, namely, when the nonparametric bias-variance tradeoff is achieved and the optimal rate of convergence of the local linear estimator is found. Also note that condition $k/n\to 0$, which is standard in nearest-neighbor estimation, follows from assuming that $k\to\infty$ and $\tau_{n,k}^{2}(x)\sqrt{k}$ is bounded.